OpenStudy (anonymous):

(2a-9)(3a-squared+4a-4) solve please. multiplying polynomials.

OpenStudy (anonymous):

no! sorry but, i know how yo do binominals, just tri nomials i need help with.

OpenStudy (whpalmer4):

@Fireflame13 Multiplying trinomials is just like multiplying binomials, there's just more work to be done. Each term in the first polynomial multiplies each term in the second. \[(2a+9)(3a^2+5a-6) = 2a(3a^2+5a-6) + 9(3a^2+5a-6) =\]\[6a^3+10a^2-12a+27a^2+45a-54 = 6a^3+37a^2+57a-54\] (I did a somewhat different problem, you get what you deserve if you just blindly copy it down as your answer)

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (whpalmer4):

By the way, the video pooja195 linked to is a good one. The second way presented to multiply polynomials is perhaps even clearer than the way you normally do multiplication of numbers! Here's a bonus: a 3 times a 4 :-) \[(a+b+c)(a + b + c + d) = (a^2 + ab + ac + ad) + (ab + b^2 + bc + bd) +\]\[ (ac + bc +c^2 + cd) =a^2+b^2+c^2+2ab + 2ac + 2bc + ad + bd + cd\]